Write down Cp/Cv ratio for diatomic gas


1
Expert's answer
2021-01-25T14:08:44-0500

Answer


Degree of freedom for monoatomic gas for=5


The ratio of specific heat is given by


"\frac{C_p}{C_v}=\gamma\\=1+\frac{2}{f}\\=1+\frac{2}{5}\\=1.4"